DIFFERENTIAL GEOMETRY/PDE SEMINAR

Tuesday, July 8, 2003 Padelford C-36 1:30 pm (Note: Special Day and Time!)

Rectifiable Scans

Bob Hardt (Rice University)

We discuss the idea of scans as a tool for extending the notion of a rectifiable current to various infinite mass generalizations. These are suitable for some Plateau-type variational problems. We obtain existence and partial regularity results for minimizers of fractional powers of the density. Scans also seems natural to understand in metric space and provide new proofs and extensions of the Ambrosio-Kirchheim work on currents in metric spaces. This is joint work with Thierry De Pauw.

For more information about this seminar, visit the DG/PDE Seminar Web page (from the Math Department home page, www.math.washington.edu, follow the link Seminars, Colloquia, and Conferences).

The University of Washington is committed to providing access, equal opportunity and reasonable accommodation in its services, programs, activities, education and employment for individuals with disabilities. To request disability accommodation contact the Disability Services Office at least ten days in advance at: 206-543-6450/V, 206-543-6452/TTY, 206-685-7264 (FAX), or dso@u.washington.edu.